# Deep Active Learning in the Open World

### v1

@@ Abstract
Machine learning models deployed in open-world scenarios often encounter unfamiliar conditions and perform poorly in unanticipated situations. As AI systems advance and find application in safety-critical domains, effectively handling out-of-distribution (OOD) data is crucial to building open-world learning systems. In this work, we introduce ALOE, a novel active learning algorithm for open-world environments designed to enhance model adaptation by incorporating new OOD classes via a two-stage approach. First, diversity sampling selects a representative set of examples, followed by energy-based OOD detection to prioritize likely unknown classes for annotation. This strategy accelerates class discovery and learning, even under constrained annotation budgets. Evaluations on three long-tailed image classification benchmarks demonstrate that ALOE outperforms traditional active learning baselines, effectively expanding known categories while balancing annotation cost. Our findings reveal a crucial tradeoff between enhancing known-class performance and discovering new classes, setting the stage for future advancements in open-world machine learning.
@@ 1 Introduction
Modern machine learning models have achieved remarkable progress by leveraging large amounts of labeled data LeCun et al. (2015); He et al. (2016); Khosla et al. (2020). Despite this success, most models are developed for closed-world settings, assuming that both training and test data originate from the same distribution. However, this assumption does not align with real-world environments, where models are inevitably encounter out-of-distribution (OOD) data with previously unseen classes Hendrycks & Gimpel (2022); Hendrycks et al. (2022); Salehi et al. (2022). Constrained by their fixed class boundaries, traditional models often struggle to generalize effectively to these novel classes, limiting their adaptability in open-world scenarios. Furthermore, obtaining human supervision for these novel examples in open-world scenarios is often time-consuming and costly, posing a significant challenge to model adaptation and improvement. Active learning, which iteratively selects the most informative examples for labeling, has emerged as a promising approach to address the expensive nature of gathering human supervision. By prioritizing examples that provide the most significant information gain, active learning can enhance the model’s learning efficiency while reducing the need for extensive human annotation. This approach is particularly valuable in open-world scenarios, where the high annotation cost and time-consuming nature of labeling make traditional supervised learning methods impractical.
@@
Despite its potential, existing active deep learning algorithms (see (Zhan et al., 2022; Zhang et al., 2024a) for overviews) have rarely been studied under open-world scenarios, particularly those involving novel classes and imbalanced data distributions. In this work, we addresses this gap by developing a novel active learning algorithm that integrates energy-based OOD detection techniques to handle the complexity of open-world environments. Our method is designed for multi-class classification tasks in the open-world, where the model encounters both known and unknown classes after deployment. Of the few works that study active learning under open-world settings, existing methods are often tailored to specific vision tasks, which result in highly specialized algorithm designs. When these methods are simplified for more generic problems like classification, they often reduce to basic uncertainty or diversity sampling techniques. In contrast, our approach in this paper offers a more comprehensive and versatile solution. By bridging OOD detection and diversity in our sampling strategy, we provide a more comprehensive sampling strategy compared to approaches that focus on only one of these aspects.
@@
We propose ALOE (Active Learning in Open-world Environments), a two-stage algorithm tailored for open-world active learning. ALOE addresses the unique challenges of class discovery in open-world settings by dynamically incorporating new OOD classes through a structured sampling strategy: In the first stage, ALOE performs diversity sampling to select a representative set of examples from the data pool. This sampling ensures broad coverage of the data distribution, capturing a range of potential new classes and concepts. By focusing on diversity, the algorithm increases the probability of identifying and learning from rare or infrequent classes that may otherwise be overlooked in a random sampling approach. The second stage leverages energy scoring function to rank examples within each cluster, prioritizing instances most likely to belong to unknown OOD classes. The energy-based OOD detection provides a unified framework that distinguishes between in-distribution (ID) and OOD examples with greater resilience to model overconfidence. By focusing annotation efforts on these high-priority examples, ALOE accelerates the discovery and learning of new classes, even when the annotation budget is limited.
@@
To empirically validate our approach, we conducted experiments on long-tail imbalanced image classification datasets. This choice of datasets is motivated by the prevalence of long-tail distributions in real-world scenarios, where rare classes or concepts are often underrepresented. Such imbalanced distributions make random sampling ineffective in discovering unknown classes, particularly those with small sample sizes. Our experimental results demonstrate the effectiveness of our approach. Compared with random sampling, on ImageNet-LT, ALOE saves 70% of annotation cost to achieve the same accuracy. In six out of six experimental settings (5.2.2 and 5.2.3), we observed that our algorithm performs the best comparing to all baseline experiments. These highlights underscore the potential of our method to significantly enhance model adaptation in open-world environments.
@@ 2.1 Open-World Learning
Open-world learning addresses the challenge of operating in dynamic environments where the model starts with a set of known classes and must detect and manage instances from unknown classes. This paradigm involves both labeled and unlabeled data, with a mixture of known and unknown classes, requiring the model to either classify examples into known categories or recognize them as novel Rizve et al. (2022); Zhu et al. (2024); Xiao et al. (2024). Unlike novel class discovery Fini et al. (2021); Zhong et al. (2021b); Han et al. ; Zhong et al. (2021a); Roy et al. (2022), which focuses solely on discovering new object categories, open-world learning must also correctly identify instances from previously known classes, making the task more complex.
Open-world semi-supervised learning Cao et al. (2022); Sun et al. (2024) generalizes semi-supervised learning by considering scenarios where the data includes both labeled and unlabeled instances from known and unknown classes. The model must learn to classify known classes and identify novel ones from unlabeled data. In contrast to traditional supervised settings, the ability to generalize to unseen classes is crucial. Open-set recognitionGe et al. (2017); Sun et al. (2020); Geng et al. (2020) shares similarities with open-world learning by allowing the model to reject novel instances during testing. However, open-world recognition extends this by requiring the model to incrementally learn and incorporate these novel classes into the set of known categoriesBoult et al. (2019); Bendale & Boult (2015). Open-world contrastive learning Sun & Li (2022) further enhances this by learning compact representation spaces that facilitate both the classification of known classes and the discovery of novel ones.
Open-world learning represents a significant advancement due to its capacity for active instance selection. In this setting, the model actively selects instances from unknown classes for labeling, allowing it to continuously expand the set of known classes as new data is labeled. Earlier work by Bendale and Boult Bendale & Boult (2016) laid the foundation for open-world recognition, introducing the Nearest Non-Outlier (NNO) algorithm and establishing evaluation protocols for managing novel classes in a continual learning environment. More recently, a comprehensive review by Zhu et al. Zhu et al. (2024) highlighted three core components for open-world systems: rejecting unknowns, discovering novel classes, and incrementally learning from them. These components form the basis of many modern approaches to open-world learning. Our research extends this by integrating open-world learning with long-tail distributions and active sample selection, enabling the model to not only handle class imbalance but also dynamically evolve as new classes emerge.
@@ 2.2 Out-of-Distribution Detection
Out-of-distribution (OOD) detection is a fundamental challenge for machine learning models deployed in open-world environments. It involves identifying whether inputs belong to unknown classes that the model has not encountered during training. The overconfidence of neural networks when handling out-of-distribution data was first revealed in Nguyen et al. (2015). Research in OOD detection has taken several main directions: post-hoc methods that devise scoring functions for detecting OOD inputs Liu et al. (2020); Lee et al. (2018); Sun et al. (2022), training-time regularization methods that leverage additional auxiliary OOD datasets to address OOD detection Hendrycks et al. (2018); Van Amersfoort et al. (2020); Katz-Samuels et al. (2022); Bai et al. (2023; 2024), and approaches exploring representation learning, such as exploring multiview contrastive losses Khosla et al. (2020); Chen et al. (2020) for OOD detection Winkens et al. (2020); Sehwag et al. (2021); Ming et al. (2022).
In particular, post-hoc methods address OOD detection by deriving test-time OOD scoring functions for a pre-trained classifier. These methods include maximum softmax probability Hendrycks & Gimpel (2016), distance-based scores Lee et al. (2018); Sun et al. (2022), energy-based scores Liu et al. (2020), activation rectification Sun et al. (2021), ViM score Wang et al. (2022), gradient-based scores Huang et al. (2021), among others. The energy score Liu et al. (2020) provides a unified framework that more effectively distinguishes between in-distribution (ID) and OOD examples. This approach theoretically aligns with the probability density of the inputs and is less susceptible to the overconfidence issue. In this work, we employ the free energy score for input examples to identify OOD data and cluster patterns, with the aim of effectively discovering and learning new classes from the data distribution.
@@ 2.3 Deep Active Learning
Uncertainty based strategies chooses examples that are determined to be the most uncertain for the model trained on the labeled set thus far. This includes many of the traditional uncertainty metrics such as margin, entropy and confidence scores (Lewis & Gale, 1994; Tong & Koller, 2001; Balcan et al., 2006; Settles, 2009; Kremer et al., 2014). More advanced approaches for deep learning also includes Bayesian uncertainty estimation and adversarial training (Gal et al., 2017; Ducoffe & Precioso, 2018; Beluch et al., 2018). Diversity based strategies aim to choose a set of unlabeled examples that are maximally different in an embedding space. This is usually achieved by clustering and covering techniques or greedy optimization of some global submodular diversity objective (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Bıyık et al., 2019; Citovsky et al., 2021; Bhatt et al., 2024). In this paper, we explore a wide range of these diversity-based methods in conjunction with OOD detection to discover new classes. Lastly, existing approaches also take a combination of uncertainty and diversity metrics, often predicting the expected model change if an example is labeled (Ash et al., 2019; 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022).
Active learning has recently focused much attention on the closely related field of open set learning (Ning et al., 2022; Zhang et al., 2022; Bai et al., 2024; Safaei et al., 2024; Yang et al., 2024; Mao et al., 2024). In open set learning, the primary objective is to classify all examples of unknown concepts into a single out-of-distribution class. Our study, however, addresses the more challenging open-world scenario. In this setting, we aim for the learner to not only identify unknown concepts but also to further learn and classify these examples into their appropriate concept categories.
@@ 4.1 ALOE: Our Actively Learning Algorithm Under Open-world Environments
Figure 2: Illustration of our algorithm ALOE 1. Starting with a few labeled examples, the algorithm clusters all examples, ideally by their underlying classes. Each example’s OOD score is calculated. In the bottom right plot, larger dots mean larger OOD score. Clusters with higher OOD ratios are prioritized for sampling to identify new classes. Labeled examples are then added to the training set, and the process iterates, expanding the labeled pool with each batch.
The algorithm proceeds iteratively for $T$ iterations, with each iteration $t$ comprising several key steps, as illustrated in Figure 2. Initially, a deep neural network undergoes is trained on the current labeled set $L_{t-1}$, resulting in an updated model $f^{t-1}$ that incorporates information from the newly labeled data. Subsequently, to ensure diversity in the selection process, the algorithm embeds the pool of unlabeled set $X^{\prime}$ using the neural network features learned from $f^{t-1}$. The clustering method, k-means, is then applied to these embedded features. The number of clusters $2\cdot\max(B,|\mathcal{K}_{t}|)$ is set so that we obtain a surplus of clusters to effectively filter out the in-distribution examples, where $B$ is the batch size and $|\mathcal{K}_{t}|$ is the number of annotated classes in step $t$. Empirically throughout our experiments, we find the multiplier value $2$ to be a good and robust value. It can be imagined that other clustering methods can be used. While it will be discussed more in our experiments 5.2.5, we stick to k-means when describing ALOE.
@@ 5.2.1 OOD Scoring Functions and Baseline Active Learning Algorithms
Baseline Active Learning Algorithms: We evaluated our algorithm against several standard active learning baselines: Random, Margin, Coreset, Galaxy, and Badge. Random is a straightforward baseline where examples are selected uniformly at random from the pool of unlabeled data, offering a comparison with non-strategic sampling. Margin (Scheffer et al., 2001) is an uncertainty-based sampling method that selects examples based on the difference between the top two predicted class probabilities. This approach aims to identify samples near decision boundaries, where the model is most uncertain. Coreset (Sener & Savarese, 2017) is a diversity-based approach that seeks a representative subset of the unlabeled pool by minimizing the maximum distance between any point and the nearest selected point. Badge (Ash et al., 2019) combines uncertainty and diversity by selecting examples based on gradient embeddings to cover informative examples. Galaxy (Zhang et al., 2022) is specifically designed for imbalanced scenarios, using a balanced sampling strategy to improve performance on classes with fewer labeled examples, making it well-suited for imbalanced datasets.

### v2

@@ Abstract
Machine learning models deployed in open-world scenarios often encounter unfamiliar conditions and perform poorly in unanticipated situations. As AI systems advance and find application in safety-critical domains, effectively handling out-of-distribution (OOD) data is crucial to building open-world learning systems. In this work, we introduce ALOE, a novel active learning algorithm for open-world environments designed to enhance model adaptation by incorporating new OOD classes via a two-stage approach. First, diversity sampling selects a representative set of examples, followed by OOD detection with GradNorm scores to prioritize likely unknown classes for annotation. This strategy accelerates class discovery and learning, even under constrained annotation budgets. Evaluations on three long-tailed image classification benchmarks demonstrate that ALOE outperforms traditional active learning baselines, effectively expanding known categories while balancing annotation cost. Our findings reveal a crucial tradeoff between enhancing known-class performance and discovering new classes, setting the stage for future advancements in open-world machine learning. Code of the work will be be uploaded at https://github.com/EfficientTraining/LabelBench.
@@ 1 Introduction
Modern machine learning models have achieved remarkable progress by leveraging large amounts of labeled data (LeCun et al., 2015; He et al., 2016; Khosla et al., 2020). Despite this success, most models are developed for closed-world settings, assuming that both training and test data originate from the same distribution. However, this assumption does not align with real-world environments, where models are inevitably encounter out-of-distribution (OOD) data with previously unseen classes (Hendrycks & Gimpel, 2022; Hendrycks et al., 2022; Salehi et al., 2022). Constrained by their fixed class boundaries, traditional models often struggle to generalize effectively to these novel classes, limiting their adaptability in open-world scenarios. Furthermore, obtaining human supervision for these novel examples in open-world scenarios is often time-consuming and costly, posing a significant challenge to model adaptation and improvement. Active learning, which iteratively selects the most informative examples for labeling, has emerged as a promising approach to address the expensive nature of gathering human supervision. By prioritizing examples that provide the most significant information gain, active learning can enhance the model’s learning efficiency while reducing the need for extensive human annotation. This approach is particularly valuable in open-world scenarios, where the high annotation cost and time-consuming nature of labeling make traditional supervised learning methods impractical.
@@
Despite its potential, existing active deep learning algorithms (see Zhan et al. (2022); Zhang et al. (2024a) for overviews) have rarely been studied under open-world scenarios, particularly those involving novel classes and imbalanced data distributions. In this work, we addresses this gap by developing a novel active learning algorithm that integrates OOD detection techniques with GradNorm scores to handle the complexity of open-world environments. Our method is designed for multi-class classification tasks in the open-world, where the model encounters both known and unknown classes after deployment. Of the few works that study active learning under open-world settings, existing methods are often tailored to specific vision tasks, which result in highly specialized algorithm designs. When these methods are simplified for more generic problems like classification, they often reduce to basic uncertainty or diversity sampling techniques. In contrast, our approach in this paper offers a more comprehensive and versatile solution. By bridging OOD detection and diversity in our sampling strategy, we provide a more comprehensive sampling strategy compared to approaches that focus on only one of these aspects.
@@
We propose ALOE (Active Learning in Open-world Environments), a two-stage algorithm tailored for open-world active learning. ALOE addresses the unique challenges of class discovery in open-world settings by dynamically incorporating new OOD classes through a structured sampling strategy: In the first stage, ALOE performs diversity sampling to select a representative set of examples from the data pool. This sampling ensures broad coverage of the data distribution, capturing a range of potential new classes and concepts. By focusing on diversity, the algorithm increases the probability of identifying and learning from rare or infrequent classes that may otherwise be overlooked in a random sampling approach. The second stage leverages GradNorm scoring function to rank examples within each cluster, prioritizing instances most likely to belong to unknown OOD classes. The OOD detection with GradNorm scores provides a unified framework that distinguishes between in-distribution (ID) and OOD examples with greater resilience to model overconfidence. By focusing annotation efforts on these high-priority examples, ALOE accelerates the discovery and learning of new classes, even when the annotation budget is limited.
@@
To empirically validate our approach, we conducted experiments on long-tail imbalanced image classification datasets. This choice of datasets is motivated by the prevalence of long-tail distributions in real-world scenarios, where rare classes or concepts are often underrepresented. Such imbalanced distributions make random sampling ineffective in discovering unknown classes, particularly those with small sample sizes. Our experimental results demonstrate the effectiveness of our approach. Compared with random sampling, on ImageNet-LT, ALOE saves 70% of annotation cost to achieve the same accuracy. In six out of six experimental settings (Sections 5.2.2 and 5.2.3), we observed that our algorithm performs the best comparing to all baseline experiments. These highlights underscore the potential of our method to significantly enhance model adaptation in open-world environments.
@@ 2.1 Open-World Learning
Open-world learning addresses the challenge of operating in dynamic environments where the model starts with a set of known classes and must detect and manage instances from unknown classes. This paradigm involves both labeled and unlabeled data, with a mixture of known and unknown classes, requiring the model to either classify examples into known categories or recognize them as novel (Rizve et al., 2022; Zhu et al., 2024; Xiao et al., 2024). Unlike novel class discovery that focuses solely on discovering new object categories (Fini et al., 2021; Zhong et al., 2021b; Han et al., ; Zhong et al., 2021a; Roy et al., 2022), open-world learning must also correctly identify instances from previously known classes, making the task more complex.
Open-world semi-supervised learning (Cao et al., 2022; Sun et al., 2024) generalizes semi-supervised learning by considering scenarios where the data includes both labeled and unlabeled instances from known and unknown classes. The model must learn to classify known classes and identify novel ones from unlabeled data. In contrast to traditional supervised settings, the ability to generalize to unseen classes is crucial. Open-set recognition shares similarities with open-world learning by allowing the model to reject novel instances during testing (Ge et al., 2017; Sun et al., 2020; Geng et al., 2020). However, open-world recognition extends this by requiring the model to incrementally learn and incorporate these novel classes into the set of known categories (Boult et al., 2019; Bendale & Boult, 2015). Open-world contrastive learning further enhances this by learning compact representation spaces that facilitate both the classification of known classes and the discovery of novel ones (Sun & Li, 2022).
Open-world learning represents a significant advancement due to its capacity for active instance selection. In this setting, the model actively selects instances from unknown classes for labeling, allowing it to continuously expand the set of known classes as new data is labeled. Earlier work by Bendale & Boult (2016) laid the foundation for open-world recognition, introducing the Nearest Non-Outlier (NNO) algorithm and establishing evaluation protocols for managing novel classes in a continual learning environment. More recently, a comprehensive review by Zhu et al. (2024) highlighted three core components for open-world systems: rejecting unknowns, discovering novel classes, and incrementally learning from them. These components form the basis of many modern approaches to open-world learning. Our research extends this by integrating open-world learning with long-tail distributions and active sample selection, enabling the model to not only handle class imbalance but also dynamically evolve as new classes emerge.
@@ 2.2 Out-of-Distribution Detection
Out-of-distribution (OOD) detection is a fundamental challenge for machine learning models deployed in open-world environments. It involves identifying whether inputs belong to unknown classes that the model has not encountered during training. The overconfidence of neural networks when handling out-of-distribution data was first revealed in Nguyen et al. (2015). Research in OOD detection has taken several main directions: post-hoc methods that devise scoring functions for detecting OOD inputs (Liu et al., 2020; Lee et al., 2018; Sun et al., 2022), training-time regularization methods that leverage additional auxiliary OOD datasets to address OOD detection (Hendrycks et al., 2018; Van Amersfoort et al., 2020; Katz-Samuels et al., 2022; Bai et al., 2023; 2024), and approaches exploring representation learning, such as exploring multiview contrastive losses (Khosla et al., 2020; Chen et al., 2020) for OOD detection (Winkens et al., 2020; Sehwag et al., 2021; Ming et al., 2022).
In particular, post-hoc methods address OOD detection by deriving test-time OOD scoring functions for a pre-trained classifier. These methods include maximum softmax probability (Hendrycks & Gimpel, 2016), distance-based scores (Lee et al., 2018; Sun et al., 2022), energy-based scores (Liu et al., 2020), activation rectification (Sun et al., 2021), ViM score (Wang et al., 2022a), gradient-based scores (Huang et al., 2021), among others. The GradNorm (Huang et al., 2021) score considers the gradient information providing a direct measure of how uncertain the model is with respect to its parameters. In this work, we employ the GradNorm score for input examples to identify OOD data and cluster patterns, with the aim of effectively discovering and learning new classes from the data distribution.
@@ 2.3 Deep Active Learning
Uncertainty based strategies chooses examples that are determined to be the most uncertain for the model trained on the labeled set thus far. This includes many of the traditional uncertainty metrics such as margin, entropy and confidence scores (Lewis & Gale, 1994; Tong & Koller, 2001; Balcan et al., 2006; Settles, 2009; Kremer et al., 2014; Wang et al., 2022b). More advanced approaches for deep learning also includes Bayesian uncertainty estimation and adversarial training (Gal et al., 2017; Ducoffe & Precioso, 2018; Beluch et al., 2018). Diversity based strategies aim to choose a set of unlabeled examples that are maximally different in an embedding space. This is usually achieved by clustering and covering techniques or greedy optimization of some global submodular diversity objective (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Bıyık et al., 2019; Citovsky et al., 2021; Bhatt et al., 2024). In this paper, we explore a wide range of these diversity-based methods in conjunction with OOD detection to discover new classes. Lastly, existing approaches also take a combination of uncertainty and diversity metrics, often predicting the expected model change if an example is labeled (Ash et al., 2019; 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022).
Active learning has recently focused much attention on the closely related field of open set learning (Karamcheti et al., 2021; Ning et al., 2022; Zhang et al., 2022; Bai et al., 2024; Safaei et al., 2024; Yang et al., 2024; Mao et al., 2024). In open set learning, the primary objective is to classify all examples of unknown concepts into a single out-of-distribution class. Our study, however, addresses the more challenging open-world scenario. In this setting, we aim for the learner to not only identify unknown concepts but also to further learn and classify these examples into their appropriate concept categories.
@@ 4.1 ALOE: Our Actively Learning Algorithm Under Open-world Environments
Figure 2: Illustration of our algorithm ALOE in Algorithm 1. Starting with a few labeled examples, the algorithm clusters all examples, ideally by their underlying classes. Each example’s OOD score is calculated. In the bottom right plot, larger dots mean larger OOD score. Clusters with higher OOD ratios are prioritized for sampling to identify new classes. Labeled examples are then added to the training set, and the process iterates, expanding the labeled pool with each batch.
The algorithm proceeds iteratively for $T$ iterations, with each iteration $t$ comprising several key steps, as illustrated in Figure 2. Initially, a deep neural network undergoes is trained on the current labeled set $L_{t-1}$, resulting in an updated model $f^{t-1}$ that incorporates information from the newly labeled data. Subsequently, to ensure diversity in the selection process, the algorithm embeds the pool of unlabeled set $X^{\prime}$ using the neural network features learned from $f^{t-1}$. The clustering method, k-means, is then applied to these embedded features. The number of clusters $2\cdot\max(B,|\mathcal{K}_{t}|)$ is set so that we obtain a surplus of clusters to effectively filter out the in-distribution examples, where $B$ is the batch size and $|\mathcal{K}_{t}|$ is the number of annotated classes in step $t$. Empirically throughout our experiments, we find the multiplier value $2$ to be a good and robust value. It can be imagined that other clustering methods can be used. While it will be discussed more in our experiments in Section 5.2.5, we stick to k-means when describing ALOE.
@@ 5.1 Experiment Setup
In our experiment, we utilize the cold starting approach by reinitializing the model from the CLIP model checkpoint after every iteration of annotation. Specifically, for labeled set $L_{t}$ with $|\mathcal{K}^{t}|$ known classes, we attach a linear head on the CLIP image encoder model with $|\mathcal{K}^{t}|$ outputs. We then finetune this model on the labeled set $L_{t}$ for each iteration $t$.
@@ 5.2.1 OOD Scoring Functions and Baseline Active Learning Algorithms
Baseline Active Learning Algorithms: We evaluated our algorithm against several standard active learning baselines: Random, Margin, Coreset, Galaxy, Badge, ProbCover, TypiClust, and EOAL. Random is a straightforward baseline where examples are selected uniformly at random from the pool of unlabeled data, offering a comparison with non-strategic sampling. Margin (Scheffer et al., 2001) is an uncertainty-based sampling method that selects examples based on the difference between the top two predicted class probabilities. This approach aims to identify samples near decision boundaries, where the model is most uncertain. Coreset (Sener & Savarese, 2017) is a diversity-based approach that seeks a representative subset of the unlabeled pool by minimizing the maximum Euclidean distance between any point in the dataset and the nearest selected point in the subset. Badge (Ash et al., 2019) combines uncertainty and diversity by selecting examples based on gradient embeddings to cover informative examples. Galaxy (Zhang et al., 2022) is specifically designed for imbalanced scenarios, using a balanced sampling strategy to improve performance on classes with fewer labeled examples, making it well-suited for imbalanced datasets. ProbCover (Yehuda et al., 2022) adopts a covering-based approach to maximize probabilistic coverage over the data distribution, effectively identifying critical examples under a constrained labeling budget. It is important to note that ProbCover is highly sensitive to its hyperparameter. As active learning can only run onnce when deployed in practice, we selected a single value that is reasonably suitable across various scenarios. TypiClust (Hacohen et al., 2022) is a low-budget active learning method that prioritizes selecting samples from dense regions of the feature space, ensuring representative coverage even with limited labeling resources. Entropic Open-set Active Learning (EOAL(Safaei et al., 2024) addresses open-set scenarios by employing entropy-based selection to handle both known and unknown classes effectively, ensuring robust performance in open-world settings. The time cost analysis of all baselines and ALOE is included in Appendix C.